OpenStudy (anonymous):

Can someone please help me I am stuck... How do the graphs of y=1/x and y=5/x+6 compare?

OpenStudy (stamp):

\[y_{1}=\frac{1}{x},\ y_{2}=\frac{5}{x+6}\]Can you visualize y1 or y2? If so, draw a sketch of the graph(s)

OpenStudy (stamp):

It might help to think of y2 as\[y_{2}=5(\frac{1}{x+6})\]

OpenStudy (anonymous):

What happened to the other x?